OpenStudy (anonymous):

Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <6, -2>, v = <2, 6>

OpenStudy (anonymous):

Two vectors are orthogonal if their dot product is zero. To check if parallel/neither, use the dot product formula: \[u\cdot v=\|u\|~\|v\|~\cos\theta\] solve for \(\theta\). (\(\|u\|\) is the magnitude of \(u\).)

OpenStudy (anonymous):

i got orthogonal

OpenStudy (anonymous):

actuallu neither

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

actuallu neither

OpenStudy (anonymous):

No, you were right the first time! \[u\cdot v=(6\times2)+(-2\times6)=12-12=0\] So they're orthogonal.

OpenStudy (anonymous):

thank you!!!